# Weighted Fisher divergence for high-dimensional Gaussian variational inference

### v1

@@ 6.1 Interpretation and related methods
Previously, Elkhalil et al. (2021) have designed autoencoders based on minimizing a batch approximation of the Fisher divergence using SGD. Cai et al. (2024) also proposed a BaM algorithm that derived closed form updates of $(\mu,\Sigma)$ by minimizing the objective,
@@
with respect to $\lambda$ at the $t$th iteration, where $\rho_{t}=Bd/t$ is the learning rate. BaM can be interpreted as a proximal point method that produces a sequence of variational densities $q_{0},q_{1},\dots$ such that $q_{t+1}$ matches the scores $g_{h}(\theta)$ at $\{\theta_{1},\dots,\theta_{B}\}$ on average better than $q_{t}$, while the KLD based penalty term ensures stability by preventing $q_{t+1}$ from deviating too much from $q_{t}$. Similarly, the SGD algorithm in Table 5 can be interpreted as minimizing
@@
which corresponds to the SGD update with step size $\rho_{t}$ employed in Table 5. This discussion extends similarly to the Fisher divergence.
@@
Instead of viewing (9) as biased estimates of the gradients of SD and FD, we can consider $\hat{S}_{q_{t}}(\lambda)$ and $\hat{F}_{q_{t}}(\lambda)$ as new objectives, which measure the divergence between $q_{\lambda}(\theta)$ and $p(\theta|y)$ based on their gradients evaluated at randomly selected samples at each iteration. Indeed, $\hat{S}_{q_{t}}(\lambda)$ and $\hat{F}_{q_{t}}(\lambda)$ reduce to zero when $q_{\lambda}(\theta)=p(\theta|y)$. At each iteration $t$, $q_{t+1}$ updates $q_{t}$ so as to reduce the difference in gradients between $q_{t}$ and the true posterior when evaluated on the randomly selected batch of samples. As $q_{t}$ converges to $p(\theta|y)$, the samples also shift towards the region where the true posterior has high probability mass. Figure 5 illustrates the progression of SDb with batch size $B=5$ using ${\text{N}}(2,1.5)$ as target. As optimization proceeds, SDb gradually refines the variational density by emulating the gradients of the target on the batch of samples and converges steadily toward the target.
@@
The SGD algorithm in Table 5 differs from BaM in several key aspects. While BaM relies on KLD regularization to stabilize updates and has closed-form updates for $(\mu,\Sigma)$, we use an $\ell_{2}$ penalty and a linearization of the batch approximation leading to SGD. By avoiding linearization of $\hat{S}_{q_{t}}(\lambda)$ and SGD, the number of iterations required for convergence is reduced in BaM, but each BaM iteration is expensive as the closed form update of $\Sigma$ involves inverting a $d\times d$ matrix with cost of $\mathcal{O}(d^{3})$, although this can be reduced to $\mathcal{O}(d^{2}B+B^{3})$ for small batch size $B$ through low rank solvers. This high cost can result in long runtimes in high dimensions. Moreover, BaM is designed for full covariance matrices and it is not clear how sparsity can be enforced in the precision matrix to take advantage of the posterior conditional independence structure in hierarchical models. BaM can also run into instability and numerical issues with ill-conditioned matrices in practice, which may not be alleviated even with larger batch sizes. On the other hand, SGD allows updating of the Cholesky factor of the precision matrix, where sparse structures can be easily enforced. Smaller batches can also be used, which further reduces the computation and storage burden. While BaM is suited for full covariance Gaussian VI, our approach provides a more scalable and stable alternative for high-dimensional hierarchical models with conditional independence structure.
@@ 6.2 Batch approximated objective under mean-field
Next, we investigate behavior of the batch approximated FD and SD under the mean-field assumption considered in Section 2. Suppose the target $p(\theta|y)$ is ${\text{N}}(\nu,\Lambda^{-1})$ with non-diagonal precision matrix $\Lambda$, and the variational approximation $q(\theta)$ is ${\text{N}}(\mu,\Sigma)$ where $\Sigma$ is a diagonal matrix. Using $B>1$ samples $\{\theta_{1},\dots,\theta_{B}\}$ from an estimate of $q$, $\hat{q}(\theta)={\text{N}}(\hat{\mu},\hat{\Sigma})$, where $\hat{\Sigma}$ is also a diagonal matrix, the batch approximated SD and FD are
@@
Proof.
$\nabla_{\Sigma_{ii}}\hat{S}_{q}(\lambda)$ $=V_{ii}-U_{ii}\Sigma_{ii}^{-2},$
Setting these derivatives to zero yields
@@
Solving these equations simultaneously, we obtain
$\Sigma^{\hat{F}}_{ii}$ =$=-\{C_{\theta,ii}+(\Sigma_{ii}^{\hat{F}})^{2}\overline{g}_{h,i}^{2}\}/\{C_{\theta g,ii}-\Sigma_{ii}^{\hat{F}}\overline{g}_{h,i}^{2}\}$
$\quad\implies\Sigma^{\hat{F}}_{ii}=-C_{\theta,ii}/C_{\theta g,ii}.$
Plugging these values into (10) yields corresponding values for $\mu_{i}^{\hat{S}}$ and $\mu_{i}^{\hat{F}}$. ∎
Next, we study limiting behavior of the batch approximated SD and FD as the batch size $B\to\infty$. Theorem 2 relies on the limits of summary statistics in (8) presented in Lemma 3.
Lemma 3.
Suppose $\{\theta_{1},\dots,\theta_{B}\}$ are samples from $\hat{q}(\theta)={\text{N}}(\hat{\mu},\hat{\Sigma})$ and the target is $p(\theta|y)={\text{N}}(\nu,\Lambda^{-1})$. As $B\rightarrow\infty$, the summary statistics in (8) satisfy
@@
Proof.
[…]
Theorem 2.
Suppose the target $p(\theta|y)$ is ${\text{N}}(\nu,\Lambda^{-1})$. Let the variational approximation $q(\theta)$ be ${\text{N}}(\mu,\Sigma)$, and $\hat{q}(\theta)={\text{N}}(\theta|\hat{\mu},\hat{\Sigma})$ be an estimate of $q(\theta)$, where $\Sigma$ and $\hat{\Sigma}$ are both diagonal matrices. As the batch size $B\rightarrow\infty$, the batch approximated SD and FD are minimized at $(\mu^{\hat{S}},\Sigma^{\hat{S}})$ and $(\mu^{\hat{F}},\Sigma^{\hat{F}})$ respectively, where
@@
Proof.
Results can be obtained by applying the continuous mapping theorem on Lemma 2 and using the results in Lemma 3. Note that $(\Lambda\hat{\Sigma}\Lambda)_{ii}=\sum_{j=1}^{d}\hat{\Sigma}_{jj}\Lambda_{ij}^{2}$ and $(\hat{\Sigma}\Lambda)_{ii}=\hat{\Sigma}_{ii}\Lambda_{ii}$. ∎
@@
7 Applications
We evaluate the performance of algorithms in Tables 4 and 5 by applying them to logistic regression, GLMMs and stochastic volatility models, and compare their results with BaM and MCMC. MCMC sampling is performed using RStan by running two chains in parallel, each with 50,000 iterations. The first half is discarded as burn-in, while the remaining 50,000 draws are used to compute kernel density estimates, regarded as the gold standard.
$\textbf{x}_{v}^{(1)},\dots,\textbf{x}_{v}^{(m)}$ and $\textbf{x}_{g}^{(1)},\dots,\textbf{x}_{g}^{(m)}$ represent independent samples drawn from the variational approximation and MCMC respectively, $k$ is the radial basis kernel function and $m=1000$. $M^{*}$ is computed 50 times for each variational approximation and a higher value indicates better multivariate accuracy. In addition, we assess the ability to capture the marginal mode of each variable accurately using the normalized absolute difference in mode ($|\mu-\mu^{*}|/\sigma^{*}$), where $\mu^{*}$ and $\sigma^{*}$ denote the mode and standard deviation of each variable based on MCMC samples.
To assess convergence, we track unbiased estimates of the lower bound, $\hat{{\mathcal{L}}}$, averaged over every 1000 iterations for SGD methods and 50 iterations for BaM, to reduce noise. Less iterations are used for averaging in BaM, as it uses closed form updates, which leads to more stable trajectories with reduced oscillations. Moreover, BaM usually requires a larger batch size and converges faster than SGD methods. Each algorithm is terminated when the gradient of a linear regression line fitted to the past five lower bound averages is less than 0.01, or when the maximum number of iterations is reached. The batch size of FDb and SDb is adjusted according to model complexity, and larger batch sizes are used for more complex models. All experiments are performed on a 16GB Apple M1 computer, using R and Julia 1.11.2.
7.1 Logistic regression
KLD FDr SDr FDb SDb BaM MCMC
MMD German 8.82$\pm$1.72 1.11$\pm$0.02 0.75$\pm$0.01 2.86$\pm$0.07 8.37$\pm$1.28 8.29$\pm$1.35 -
a4a 2.49$\pm$0.03 0.71$\pm$0.02 0.71$\pm$0.01 1.61$\pm$0.02 3.04$\pm$0.04 5.22$\pm$0.08 -
$\frac{|\mu-\mu^{*}|}{\sigma^{*}}$ German 0.08$\pm$0.05 0.55$\pm$0.68 0.74$\pm$0.62 0.25$\pm$0.27 0.08$\pm$0.06 0.08$\pm$0.05 -
a4a 0.18$\pm$0.15 0.45$\pm$0.52 0.38$\pm$0.43 0.40$\pm$0.43 0.15$\pm$0.15 0.14$\pm$0.16 -
time German 3.2 (35) 15.0 (60) 9.8 (39) 6.2 (26) 3.8 (15) 1.6 (0.35) 354.4
a4a 19.2 (44) 108.5 (31) 55.2 (16) 14.0 (10) 70.4 (48) 13.6 (1.05) 28556.1
@@ 7.1 Logistic regression
Table 6: Logistic regression. Estimates of mean and standard deviation for MMD and normalized absolute difference in mode.
Runtime is in seconds and number of iterations (in thousands) is given in brackets.
@@
Consider a logistic regression model where $y=(y_{1},\dots,y_{n})^{\top}$ represents $n$ independent binary responses. Each $y_{i}$ follows a Bernoulli distribution with success probability $p_{i}$, modeled as
$\text{\rm logit}(p_{i})=X_{i}^{\top}\theta\quad\text{for}\quad i=1,\dots,n,$
@@
where $X_{i}\in\mathbb{R}^{d}$ denotes the covariates of the $i$th observation and $\theta\in\mathbb{R}^{d}$ denotes the unknown coefficients, which is assigned the prior ${\text{N}}(0,\sigma_{0}^{2}I_{d})$ with $\sigma_{0}^{2}=100$. Here, the precision matrix of the Gaussian variational approximation is not sparse, but a full matrix. The log joint density of the model, gradient and Hessian are given by
$\nabla_{\theta}\log h(\theta)$ $=X^{\top}(y-w)-\theta/\sigma_{0}^{2},$
$\nabla_{\theta}^{2}\log h(\theta)$ $=-X^{\top}WX-I_{d}/\sigma_{0}^{2},$
@@
We fit the logistic regression model to two real datasets from the UCI machine learning repository. The first is the German credit data, which consists of $1000$ individuals classified as having a “good” or “bad” credit risk, and $20$ attributes. All quantitative predictors are standardized to have mean zero and standard deviation one, while qualitative predictors are encoded using dummy variables. The second is the Adult data with $48,842$ observations, which is used to predict whether an individual’s annual income exceeds \$$50,000$ based on $14$ attributes. For MCMC to be feasible, we use the preprocessed a4a data at www.csie.ntu.edu.tw/~cjlin/libsvmtools/datasets/binary.html, which has $4781$ training samples derived from the Adult data. After preprocessing, $d=49$ for German credit data and $d=124$ for a4a data. As the a4a data is high dimensional and has a large number of observations, we only generate $10,000$ MCMC samples from two parallel chains, each consisting of $10,000$ iterations. We use a batch size of $B=3$ for FDb and SDb, and $B=50$ for BaM. The maximum number of iterations is $60,000$.
Figure 6: German credit and a4a data. Average lower bound over every 1000 iterations.
@@
Fig 6 shows progression of the lower bound for SGD methods. FDr and SDr converge very slowly, and attain much poorer lower bounds than other methods at termination. This is likely due to the high variance in their gradient estimates, as discussed in Section 5.1. In contrast, SDb converges rapidly and achieves the highest lower bound within the first one thousand iterations, surpassing even KLD. The lower bound achieved by FDb, on the other hand, is lower than KLD and SDb although it performs better than FDr and SDr.
@@
From the MMD results in Table 6, FDr and SDr produce much poorer variational approximations than KLD, while FDb and SDb provide significant improvements over FDr and SDr. In particular, SDb produced results comparable to KLD for the German credit data, and better than KLD for the a4a data. In terms of runtime, KLD is the fastest among SGD methods. For the German credit data, FDr, SDr, FDb and SDb each takes $\sim$0.25s per 1000 iterations, but FDb and SDb require less iterations to converge. For the a4a data, FDr and SDr require $\sim$3.5s per 1000 iterations compared to $\sim$1.5s for FDb and SDb, due to the larger cost of computing the Hessian in high dimensions. BaM converges most rapidly, outperforming all SGD methods in runtime. Its MMD values are comparable to KLD for the German credit data, and much higher than SGD methods for the a4a data.
Overall, MMD values for optimizing FD and SD based on batch approximation are consistently higher than those based on the reparameterization trick. In terms of capturing the marginal mode accurately, SDb and BaM are comparable to KLD for the German credit data, and better than KLD for the more challenging a4a data. All variational approximations are at least an order of magnitude faster than MCMC, and the speedup is more apparent as the dimension and number of observations increase.
7.2 Generalized linear mixed models

### v2

@@ 6.1 Interpretation and related methods
Previously, Elkhalil et al. (2021) designed autoencoders based on minimizing a batch approximation of the Fisher divergence using SGD. Cai et al. (2024) also proposed a BaM algorithm that derived closed form updates of $(\mu,\Sigma)$ by minimizing the objective,
@@
with respect to $\lambda$ at the $t$th iteration, where $\rho_{t}=Bd/t$ is the learning rate. BaM can be interpreted as a proximal point method that produces a sequence of variational densities $q_{0},q_{1},\dots$ such that $q_{t+1}$ matches the scores $g_{h}(\theta)$ at $\{\theta_{1},\dots,\theta_{B}\}$ on average better than $q_{t}$, while the KLD based penalty ensures stability by preventing $q_{t+1}$ from deviating too much from $q_{t}$. Similarly, Algorithm 2 can be interpreted as minimizing
@@
which corresponds to the SGD update with stepsize $\rho_{t}$ employed in Algorithm 2. This discussion extends similarly to the FD.
@@
Instead of viewing (9) as biased estimates of the gradients of SD and FD, we can consider $\hat{S}_{q_{t}}(\lambda)$ and $\hat{F}_{q_{t}}(\lambda)$ as new objectives, which measure the divergence between $q_{\lambda}(\theta)$ and $p(\theta|y)$ based on their gradients evaluated at randomly selected samples at each iteration. Indeed, $\hat{S}_{q_{t}}(\lambda)$ and $\hat{F}_{q_{t}}(\lambda)$ reduce to zero when $q_{\lambda}(\theta)=p(\theta|y)$, which can be seen from (8) by plugging in $g_{h}(\theta)=-\Sigma^{-1}(\theta-\mu)$, as each term in the sums is equal to zero. This supports their use as optimization objectives. At each iteration $t$, $q_{t+1}$ updates $q_{t}$ so as to reduce the difference in gradients between $q_{t}$ and the true posterior when evaluated on the randomly selected batch of samples. As $q_{t}$ converges to $p(\theta|y)$, the samples also shift towards the region where the true posterior has high probability mass. Figure 5 illustrates the progression of SDb with batchsize $B=5$ using ${\text{N}}(2,1.5)$ as target. As optimization proceeds, SDb gradually refines the variational density by emulating the gradients of the target on the batch of samples and converges steadily toward the target.
Despite the preceding discussion and empirical evidence in Section 7, it is important to establish formal convergence for SDb and FDb, as these algorithms can be interpreted as either relying on biased gradients of the SD and FD, or unbiased gradients based on objectives that change at each iteration, and convergence is not guaranteed in either case. In this article, we do not resolve these issues definitively, and we will leave these as open problems for future work. However, we make some contribution in this direction by providing Theorem 4 below, which proves the convergence of SDb in the limit of infinite batchsize with a Gaussian target, where natural gradients (Tan, 2025) are used with a constant stepsize. The proof of Theorem 4 given in the supplement follows that of BaM in Cai et al. (2024) closely under similar settings. There are several differences between the conditions in Theorem 4 and SDb as implemented in Algorithm 2. In particular, we update $T_{t}$ instead of $\Sigma_{t}^{-1}$, using an adaptive instead of constant stepsize, based on Euclidean rather than natural gradients, using a finite instead of infinite batchsize.
Theorem 4.
Suppose the target is ${\text{N}}(\mu,\Lambda)$ and the variational approximation at iteration $t$ is ${\text{N}}(\mu_{t},\Sigma_{t})$, where $\lambda_{t}=(\mu_{t},\Sigma_{t})$. Define the normalized errors,
$\epsilon_{t}$ $=\Lambda^{1/2}(\mu_{t}-\nu),\quad\Delta_{t}$ $=\Lambda^{-1/2}(\Sigma_{t}^{-1}-\Lambda)\Lambda^{-1/2}.$
If $\hat{S}_{q}(\lambda)$ is minimized using SGD using a constant stepsize $0<\rho<1/4$, based on the natural gradients updates,
$\Sigma_{t+1}^{-1}=\Sigma_{t}^{-1}+2\rho\nabla_{\Sigma}\hat{S}_{q_{t}}(\lambda_{t}),$
$\mu_{t+1}=\mu_{t}-\rho\Sigma_{t+1}\nabla_{\mu}\hat{S}_{q_{t}}(\lambda_{t}),$
then in the limit of infinite batch size ($B\to\infty$), $\|\epsilon_{t}\|\to 0$ and $\|\Delta_{t}\|\to 0$ as $t\to\infty$, where $\|\cdot\|$ denotes the spectral norm.
@@
Algorithm 2 differs from BaM in several key aspects. While BaM relies on KLD regularization for stability and has closed-form updates for $(\mu,\Sigma)$, we use an $\ell_{2}$ penalty and a linearization of the batch approximation leading to SGD. By avoiding linearization of $\hat{S}_{q_{t}}(\lambda)$ and SGD, the number of iterations required for convergence is reduced in BaM, but each BaM iteration is expensive as the closed form update of $\Sigma$ involves inverting a $d\times d$ matrix with cost of $\mathcal{O}(d^{3})$, although this can be reduced to $\mathcal{O}(d^{2}B+B^{3})$ for small batchsize $B$ through low rank solvers. This high cost can result in long runtimes in high dimensions. Moreover, BaM is designed for full covariance matrices and it is not clear how sparsity can be enforced in the precision matrix to take advantage of the posterior conditional independence structure in hierarchical models. BaM can also run into instability and numerical issues with ill-conditioned matrices in practice, which may not be alleviated even with larger batchsizes. On the other hand, SGD allows updating of the Cholesky factor of the precision matrix, where sparse structures can be easily enforced. Smaller batchsizes can also be used, which further reduces the computation and storage burden. While BaM is suited for full covariance Gaussian VI, our approach provides a more scalable and stable alternative for high-dimensional hierarchical models with conditional independence structure.
The batch, match and patch (pBaM) algorithm (Modi et al., 2025) extends BaM to higher dimensions through a patch step, that projects the covariance matrix into a family of low rank plus diagonal matrices (Ong et al., 2018) such that $\Sigma=\Lambda\Lambda^{\top}+\Psi$, via an expectation maximization (EM) procedure. The rank $K$ of the low-rank factor $\Lambda$ controls the trade-off between computational efficiency and approximation accuracy, and the approximation accuracy improves with a larger $K$ for a given batchsize $B$. Each EM step has a cost of $\mathcal{O}(dK^{2}+K^{3}+KBd)$ while each BaM step has a cost of $\mathcal{O}(dB^{2}+B^{3}+KBd)$. Hence, pBaM is much more scalable in high-dimensional regimes, where storing and updating dense covariance matrices is impractical. Unlike pBaM, our approach does not require the tuning of additional hyperparameters such as $K$, but is still scalable and able to exploit sparsity.
@@ 6.2 Batch approximated objective under mean-field
[…]
@@
Next, we study limiting behavior of the batch approximated SD and FD as the batchsize $B\to\infty$. Theorem 5 relies on the limits of summary statistics step 6 of Algorithm 2 presented in Lemma 3.
Lemma 3.
@@
$\begin{gathered}\overline{\theta}\xrightarrow{\text{a.s.}}\hat{\mu},\quad C_{\theta}\xrightarrow{\text{a.s.}}\hat{\Sigma},\quad\overline{g}_{h}\xrightarrow{\text{a.s.}}\Lambda(\nu-\hat{\mu}),\\ C_{g}\xrightarrow{\text{a.s.}}\Lambda\hat{\Sigma}\Lambda,\quad C_{\theta g}\xrightarrow{\text{a.s.}}-\hat{\Sigma}\Lambda.\end{gathered}$
Theorem 5.
@@
$\Sigma^{\hat{S}}_{ii}\xrightarrow{\text{a.s.}}\sqrt{\frac{\hat{\Sigma}_{ii}}{\sum_{j=1}^{d}\hat{\Sigma}_{jj}\Lambda_{ij}^{2}}},\qquad\Sigma^{\hat{F}}_{ii}\xrightarrow{\text{a.s.}}\frac{1}{\Lambda_{ii}},$
$\mu_{i}^{\hat{S}}\xrightarrow{\text{a.s.}}\hat{\mu}_{i}+\sqrt{\frac{\hat{\Sigma}_{ii}}{\sum_{j=1}^{d}\hat{\Sigma}_{jj}\Lambda_{ij}^{2}}}\sum_{j=1}^{d}\Lambda_{ij}(\nu_{j}-\hat{\mu}_{j}),$
$\mu_{i}^{\hat{F}}\xrightarrow{\text{a.s.}}\hat{\mu}_{i}+\frac{1}{\Lambda_{ii}}\sum_{j=1}^{d}\Lambda_{ij}(\nu_{j}-\hat{\mu}_{j}).$
From Lemma 3, $C_{\theta g}$ converges almost surely to $\hat{\Sigma}\Lambda$, with $i$th diagonal entry $\hat{\Sigma}_{ii}\Lambda_{ii}<0$. Thus, diagonal elements of $C_{\theta g}$ are likely negative for a sufficiently large $B$, but may be positive for a small batchsize $B$. In that case, assuming $\mu_{i}=\bar{\theta}_{i}+\Sigma_{ii}\bar{g}_{h,i}$, $\nabla_{\Sigma_{ii}}\hat{F}_{q}(\lambda)=-2\Sigma_{ii}^{-2}(C_{\theta,ii}+C_{\theta g,ii})<0$, and $\hat{F}_{q}(\lambda)$ decreases monotonically as $\Sigma_{ii}\rightarrow\infty$. Thus the batch approximated FD faces the issue of “variance explosion”. This is in stark contrast to results in Section 2 where the FD has a closed form solution. On the other hand, the batch approximated SD no longer faces the issue of “variational collapse”, and has a closed form solution for any $B>1$. As $B\rightarrow\infty$, $\Sigma^{\hat{S}}_{ii}\xrightarrow{\text{a.s.}}\sqrt{\hat{\Sigma}_{ii}/(\sum_{j=1}^{d}\hat{\Sigma}_{jj}\Lambda_{ij}^{2})}$, the limit of which is equal to that in (1) where $M=\hat{\Sigma}$ in the weighted Fisher divergence. It follows from Theorem 1 that $\Sigma^{\hat{S}}_{ii}\leq\Sigma^{\hat{F}}_{ii}=\Sigma^{\text{KL}}_{ii}$ in the limit of infinite batchsize. Thus the batch approximated SD underestimates the posterior variance more severely than the batch approximated FD, for which the posterior variance estimate matches that of the KLD, as $B\rightarrow\infty$. However, unlike the FD and SD, the true mean $\nu$ is not recovered by the batch approximated FD and SD even as $B\rightarrow\infty$, unless $\Lambda$ is a diagonal matrix.
7 Applications
We evaluate the performances of Algorithms 1 and 2 by applying them to logistic regression, GLMMs and stochastic volatility models, and compare their results with BaM, pBaM and MCMC. MCMC sampling is performed using RStan by running 2 chains in parallel, each with 20,000 iterations. The first half is discarded as burn-in, while the remaining 20,000 draws are used to compute kernel density estimates, regarded as the gold standard.
As BaM allows a full covariance matrix, while pBaM uses a more restrictive factor covariance structure and hence may have lower approximation accuracy, we use BaM whenever it is computationally feasible. pBaM is only used in high-dimensional settings, where BaM is impractical or numerically unstable. The choice of batchsize for FDb, SDb, BaM and pBaM is dependent on the method and model complexity, due to the trade-off between computational efficiency and approximation accuracy. For FDb and SDb, small batchsizes are often sufficient, as only a small step is taken at each iteration due to the reliance on noisy gradient estimates in SGD. In contrast, BaM uses closed form updates that involve matrix inversion and larger batchsizes are necessary to ensure stability and avoid ill-conditioned updates.
$\textbf{x}_{v}^{(1)},\dots,\textbf{x}_{v}^{(m)}$ and $\textbf{x}_{g}^{(1)},\dots,\textbf{x}_{g}^{(m)}$ represent independent samples drawn from the variational approximation and MCMC respectively, $k$ is the radial basis kernel function and $m=1000$. $M^{*}$ is computed 50 times for each variational approximation and a higher value indicates better multivariate accuracy. In addition, we assess the ability to capture the marginal mean, mode and standard deviation of each variable accurately using the normalized absolute difference ($|\mu-\mu^{*}|/\sigma^{*}$, $|\mu-m^{*}|/\sigma^{*}$) and standard deviation ratio $\sigma/\sigma^{*}$, where $\mu$ and $\sigma$ denote the variational mean and standard deviation, and $\mu^{*}$, $m^{*}$, $\sigma^{*}$ denote the mean, mode and standard deviation of each variable based on MCMC samples. The marginal posterior mode for each variable is reported rather than the joint posterior mode. This distinction arises in high dimensions because MCMC algorithms predominantly explore the typical set, which covers most of the probability mass and is where draws tend to lie in, but this set can lie far from the neighborhood of the global mode. This shell geometry largely vanishes after marginalizing, making the marginal mode a more stable quantity to estimate from MCMC samples (Liu and Ihler, 2013; Betancourt, 2018).
To assess convergence, we track unbiased estimates of the lower bound, $\hat{{\mathcal{L}}}$, averaged over every 1000 iterations for SGD methods and 50 iterations for (p)BaM (BaM and pBaM) to reduce noise. Fewer iterations are used for averaging in (p)BaM, as it uses closed form updates, which lead to more stable trajectories. Moreover, (p)BaM usually requires a larger batchsize and converges faster than SGD methods. Each algorithm is terminated when the gradient of a linear regression line fitted to the past five lower bound averages becomes negative (this indicates that the lower bounds have reached a maximum and begun to fluctuate around it), or when the maximum number of iterations is reached. All experiments are performed on a 16GB Apple M1 computer, using R and Julia 1.11.2.
7.1 Logistic regression
KLD FDr SDr FDb SDb BaM
@@ 7.1 Logistic regression
$\frac{\sigma}{\sigma^{*}}$ German 0.99 $\pm$ 0.02 0.92 $\pm$ 0.10 0.88 $\pm$ 0.17 0.99 $\pm$ 0.02 0.99 $\pm$ 0.02 0.99 $\pm$ 0.02
a4a 0.61 $\pm$ 0.21 0.12 $\pm$ 0.12 0.22 $\pm$ 0.31 0.54 $\pm$ 0.83 0.71 $\pm$ 0.17 0.97 $\pm$ 0.12
time German 3.6 (45) 15.0 (60) 9.8 (39) 8.5 (32) 4.7 (16) 3.4 (0.9)
a4a 19.9 (45) 108.5 (31) 55.2 (16) 14.0 (10) 72.8 (49) 13.6 (1.05)
@@
Table 5: Logistic regression. Mean and standard deviation of normalized absolute difference in mode and mean, and standard deviation ratio (best values highlighted in bold). Runtime is in seconds and number of iterations (in thousands) is given in brackets.
@@
Consider the logistic regression model where $y=(y_{1},\dots,y_{n})^{\top}$ represents $n$ independent binary responses. Each $y_{i}$ follows a Bernoulli distribution with success probability $p_{i}$, modeled as
$\text{\rm logit}(p_{i})=X_{i}^{\top}\theta\quad\text{for}\quad i=1,\dots,n.$
@@
$X_{i}\in\mathbb{R}^{d}$ denotes the covariates of the $i$th observation and $\theta\in\mathbb{R}^{d}$ denotes the unknown coefficients, which is assigned the prior ${\text{N}}(0,\sigma_{0}^{2}I_{d})$ with $\sigma_{0}^{2}=100$. Here, the precision matrix of the Gaussian variational approximation is not sparse, but a full matrix. The log joint density of the model, gradient and Hessian are given in the supplement.
@@
We fit the logistic regression model to two real datasets from the UCI machine learning repository. The first is the German credit data, which consists of $1000$ individuals classified as having a “good” or “bad” credit risk, and $20$ attributes. All quantitative predictors are standardized to have mean zero and standard deviation one, while qualitative predictors are encoded using dummy variables. The second is the Adult data with $48,842$ observations, which is used to predict whether an individual’s annual income exceeds \$$50,000$ based on $14$ attributes. For MCMC to be feasible, we use the preprocessed a4a data at www.csie.ntu.edu.tw/~cjlin/libsvmtools/datasets/binary.html, which has $4781$ training samples derived from the Adult data. After preprocessing, $d=49$ for German credit data and $d=124$ for a4a data. As the a4a data has a large number of observations, we only generate $10,000$ MCMC samples from two parallel chains, each consisting of $10,000$ iterations. For these datasets, we only use BaM as it already performs very well. We use a batchsize of $B=3$ for FDb and SDb, and $B=50$ for BaM. The maximum number of iterations is $60,000$.
Figure 6: Logistic regression. First row contains plots of the lower bound averaged over every 1000 iterations and second row contains boxplots of $M^{*}$ values.
@@
Fig 6 shows the progression of the lower bound for SGD methods. FDr and SDr converge very slowly and attain much poorer lower bounds than other methods, likely due to the high variance in their gradient estimates, as discussed in Section 5.1. In contrast, SDb converges rapidly and achieves the highest lower bound within the first 1000 iterations, surpassing even KLD. The lower bound achieved by FDb is lower than KLD and SDb although it performs better than FDr and SDr. From the $M^{*}$ results in Fig 6, FDr and SDr produce much poorer variational approximations than KLD, while FDb and SDb provide significant improvements over FDr and SDr. In particular, SDb produced better results than KLD.
@@
From Table 5, KLD is the fastest among SGD methods. For the German credit data, FDr, SDr, FDb and SDb each takes $\sim$0.25s per 1000 iterations, but FDb and SDb require fewer iterations to converge. For the a4a data, FDr and SDr require $\sim$3.5s per 1000 iterations compared to $\sim$1.5s for FDb and SDb, due to the higher cost of computing the Hessian in high dimensions. BaM converges most rapidly, outperforming all SGD methods in runtime. Its $M^{*}$ values are comparable to KLD for German credit data, and much higher than SGD methods for a4a data.
Overall, $M^{*}$ values for optimizing FD and SD based on batch approximation are consistently higher than those based on the reparameterization trick. SDb and BaM are better than KLD at estimating the marginal mode and mean accurately for both datasets.
7.2 Generalized linear mixed model